One angle of a hexagon is x degree. The each of other angles are 116 degree. Find the value of x.
Answer by Alan3354(69443)

One angle of a hexagon is x degree. The each of other angles are 116 degree. Find the value of x.
------------------
There are 6 angles.
The sum of them is 720 degs
5*116 = 580
x = 720 - 580 = 140